What is the angular momentum of the hour hand of a clock If it has a length of 25 cm and a mass of 20 g?


1
Expert's answer
2021-04-30T11:24:32-0400

"L= I*\omega,\text{where:}"

"I=\frac {ml^2}{3}-\text{Inertia of a thin bar the axis of rotation} \newline
\text{passes through the end of the bar}\newline
l=25cm =0.25m;\newline
m=20g=0.02kg\newline
I=\frac{0.02*0.25^2}{3}=4,16*10^{-4}"

"\omega-\text{angular velocity}\newline
\text{one turn is }2\pi\ rad"

"\omega=\frac{2\pi}{t}"

"t=1h =3600s"

"\omega=\frac{2\pi}{3600}=1.75*10^{-3}"

"L=4.16*10^{-4}*1.75*10^{-3}=7.28*10^{-7}\newline
\text{Answer: }7.28*10^{-7}\ \frac{kg*m^2}{s}"